# The Generation Gap: Exploring Age Bias Underlying in the Value Systems of Large Language Models

### v2

@@ Appendix A World Value Survey
The WVS [^3] https://www.worldvaluessurvey.org/wvs.jsp survey is conducted every five years, which systematically probes individuals globally on social, political, economic, religious, and cultural values. We share a page of WVS questionnaire in Tab 6. See the statistics of inquiries in Fig 2. Demographic statistics of WVS is accessible via Document-Online analysis. Note that we remove ten of them that requires demographic information, as these are impossible for applying to an LLM lacking demographic data, and keep 249 inquiries as our final choices for prompting.
@@
Table 3: Prompt Pipeline Details
@@
returning null value { ”Q1”: null} map null into missing code -2
@@
redundant texts ”Answer = {‘Q1’, 1}” extract the json result
@@ Appendix B The Instability of LLM Outputs Due to Prompt Variations
    Despite potential variations in answers induced by prompt variation, we found for 95.5% of inquiries, more than half of the responses are centered on the same choice or its adjacent options. Adjacent option is a score equal to the majority score +/- 1.
@@
Appendix C Reasons of Applying PCA
  1. 1.
    Each question in WVS ought not to be equally important. Furthermore, for the questions belonging to a certain category, they correlate with each other. To this end, we need to find out the principle components among multi inquiries.
  2. 2.
    PCA here is also used as an unsupervised representation learning method. Compared to utilizing original data, the representations learned from hundreds of comparable examples (372 value vectors from different country and age groups) will mitigate the curse of dimensionality and other undesired properties of high-dimensional spaces. Other representation learning methods are also applicable. As the medium number of original dimensionality for all categories is 11, PCA is enough to handle the learning problem.
Appendix D Prompting Details
@@
The cost of API calling from Closed-coursed LLMs is less than 5 dollars. For the deployment of open-sourced models, we ran either model on a single A40 GPU with float16 precision. When prompting, we prompt models with a temperature 1.0, max token length 1024, random seed 42.
Appendix E Results on Other LLMs
In the section, we supplement the alignment ranking results on InstructGPT (Fig 7), FLAN-T5-XXL (Fig 8) and FLAN-UL2 (Fig 9) respectively.
Appendix F Results on Other Countries
We have extended our analysis to include alignment results from an additional four pairs of countries: Argentina and Brazil (Fig 7), Ethiopia and Nigeria (Fig 8), Germany and Great Britain (Fig 9), and Indonesia and Malaysia (Fig 10).
Appendix G Significant Test
In this section, we conduct two kinds of significant tests to support our study: (1) we use MANOVA to test the significant difference among human values from different age groups, and (2) we use t-distribution to test the significant tendency of LLMs towards younger group. Notes our focus lies in characterizing the inclination of LLM values towards specific age groups. So, we are claiming significant tendency instead of claiming LLMs significantly assemble of any specific age groups.
G.1 Significant Test for the Discrepancy among Human Age Groups
Our analysis should base on a reasonable precondition that in WVS, human values significantly diverse cross different age groups. We used MANOVA (multivariate analysis of variance) to test the significant difference of human values cross all age groups, as shown below:
@@ G.1 Significant Test for the Discrepancy among Human Age Groups
Result: See Tab 5. In conclusion: We reject the null hypothesis with p-value ¡ 1e-4
Country Value Num DF Den DF F Value Pr ¿ F (p-value)
@@
Table 5: P-values of value difference among different age groups in specific countries. * indicates p-value¡1e-4
G.2 Significant Test for Trend Coefficient
@@
Table 6: P-values of trend coefficients for each model on each value category. * indicates p-value¡0.1
Figure 7: Alignment rank of values of InstructGPT over different age groups in the US. Rank 1 on an specific age group represents that this age group has the narrowest gap with InstructGPT in values. A increasing monoticity indicates a closer alignment towards younger groups, vice versa.
Figure 8: Alignment rank of values of FLAN-T5-XXL over different age groups in the US. Rank 1 on an specific age group represents that this age group has the narrowest gap with FLAN-T5-XXL in values. A increasing monoticity indicates a closer alignment towards younger groups, vice versa.
Figure 9: Alignment rank of values of FLAN-UL2 over different age groups in the US. Rank 1 on an specific age group represents that this age group has the narrowest gap with FLAN-UL2 in values. A increasing monoticity indicates a closer alignment towards younger groups, vice versa.
Table 7: Alignment rank of LLMs over different age groups in Argentina and Brazil. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
Table 8: Alignment rank of LLMs over different age groups in Ethiopia and Nigeria. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
Table 9: Alignment rank of LLMs over different age groups in Gemany and Great Britain. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
Table 10: Alignment rank of LLMs over different age groups in Indonesia and Malaysia. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.

### v3

@@ Appendix A World Value Survey
The WVS333 The data can be downloaded via https://www.worldvaluessurvey.org/wvs.jsp survey is conducted every five years, which systematically probes individuals globally on social, political, economic, religious, and cultural values. We share a page of WVS questionnaire in Tab 7. See the statistics of inquiries in Fig 2. Demographic statistics of WVS are accessible via Document-Online analysis. Note that we removed ten of them that require demographic information, as these are impossible to apply to an LLM lacking demographic data, and kept 249 inquiries as our final choices for prompting.
@@
Figure 8: Prompt Pipeline Details
@@
returning null value { "Q1": null} map null into missing code -2
@@
redundant texts "Answer = {‘Q1’, 1}" extract the json result
@@ Appendix B The Instability of LLM Outputs Due to Prompt Variations
    Despite potential variations in answers induced by prompt variation, we found for 95.5% of inquiries, more than half of the responses are centered on the same choice or its adjacent options. The adjacent option is a score equal to the majority score +/- 1.
@@
Appendix C Prompting Details
@@
The cost of API calling from Closed-coursed LLMs is less than 5 dollars. For the deployment of open-sourced models, we ran either model on a single A40 GPU with float16 precision. When prompting, we prompt models with a temperature 1.0, max token length 1024, and random seed 42.
Appendix D Results on Other LLMs
In the section, we supplement the alignment ranking results on InstructGPT (Fig 9), FLAN-T5-XXL (Fig 10) and FLAN-UL2 (Fig 11), Mistral (Fig 12) and Vicuna (Fig 13) respectively.
Appendix E Results on Other Countries
We have extended our analysis to include alignment results from an additional four pairs of countries: Argentina and Brazil (Tab 14), Ethiopia and Nigeria (Tab 15), Germany and Great Britain (Tab 16), and Indonesia and Malaysia (Tab 17).
Appendix F Significance Test
In this section, we conduct two kinds of significance tests to support our study: (1) we use MANOVA to test the significant difference among human values from different age groups, and (2) we use t-distribution to test the significant tendency of LLMs towards younger groups. Notes our focus lies in characterizing the inclination of LLM values toward specific age groups. That is to say, we are claiming a significant tendency over age, rather than claiming LLMs significantly resemble any specific age group. We make a deeper discussion about our declaration in the section on Limitations.
F.1 Significance Test for the Discrepancy among Human Age Groups
Our analysis should be based on a reasonable precondition that in WVS, human values are significantly diverse across different age groups. We used MANOVA (multivariate analysis of variance) to test the significant difference in human values across all age groups, as shown below:
@@ F.1 Significance Test for the Discrepancy among Human Age Groups
Result: See Tab 4. In conclusion: We reject the null hypothesis with p-value < 1e-4
Country Value Num DF Den DF F Value Pr > F (p-value)
@@
Table 4: P-values of value difference among different age groups in specific countries. * indicates p-value<1e-4
F.2 Significance Test for Trend Coefficient
@@
Appendix G Our Consideration on Measure Design
G.1 Reasons for Applying PCA
We choose PCA for the following reasons:
  1. 1.
    Each question in WVS ought not to be equally important. Furthermore, for the questions belonging to a certain category, they correlate with each other. To this end, we need to find out the principal components among multiple inquiries.
  2. 2.
    PCA here is also used as an unsupervised representation learning method. Compared to utilizing original data, the representations learned from hundreds of comparable examples (372 value vectors from different countries and age groups) will mitigate the curse of dimensionality and other undesired properties of high-dimensional spaces. Other representation learning methods are also applicable. As the medium number of original dimensionality for all categories is 11, PCA is enough to handle the learning problem.
Furthermore, we set the target number of PCA components to three. We empirically set so, considering the medium number of original dimensionality for all categories is eleven. Then we validate this parameter by calculating the percentage of variance explained by each of the selected components. If all components are stored, the sum of the ratios is equal to 1.0. The explained variance ratio of keeping three dimensions is an average of no less than 0.72 in all categories of six models, which we believe is acceptable.
G.2 Consideration of Using the Rank of Difference as Measurement
In Sec 3.3, we utilize the rank of difference to characterize the value discrepancies and the trend coefficient over age. Presenting rank is simple and convenient for data visualization. However, using the rank of difference may ignore the magnitude (the absolute value) of difference that is (1) among the different age groups of humans or (2) between LLM values and specific age groups of humans. We further clarify that:
(1) Appx F.1 has shown significant value discrepancies among different age groups of humans in the countries we experiment on. So, using the rank of difference would not exaggerate a significant disparity between human age groups to observe, as the discrepancies have existed significantly.
(2) As shown in the second sub-figure of Fig 3, it is possible that LLMs values are far away from all human age groups. Such discrepancies also would not reflect on the rank of difference. However, our study focus lies in characterizing the inclination of LLM values towards specific age groups. That is to say, we are claiming a significant tendency over age, rather than claiming LLMs significantly resemble any specific age group. We make a deeper discussion about our declaration in the section of Limitations.
@@ G.2 Consideration of Using the Rank of Difference as Measurement
Table 5: P-values of trend coefficients for each model on each value category. * indicates p-value<0.1
Figure 9: Alignment rank of values of InstructGPT over different age groups in the US. Rank 1 on a specific age group represents that this age group has the narrowest gap with InstructGPT in values. An increasing monoticity indicates a closer alignment towards younger groups, vice versa.
Figure 10: Alignment rank of values of FLAN-T5-XXL over different age groups in the US. Rank 1 on a specific age group represents that this age group has the narrowest gap with FLAN-T5-XXL in values. An increasing monoticity indicates a closer alignment towards younger groups, vice versa.
Figure 11: Alignment rank of values of FLAN-UL2 over different age groups in the US. Rank 1 on a specific age group represents that this age group has the narrowest gap with FLAN-UL2 in values. An increasing monoticity indicates a closer alignment towards younger groups, vice versa.
Figure 12: Alignment rank of values of Mistral over different age groups in the US. Rank 1 on a specific age group represents that this age group has the narrowest gap with Mistral in values. An increasing monoticity indicates a closer alignment towards younger groups, vice versa.
Figure 13: Alignment rank of values of Vicuna over different age groups in the US. Rank 1 on a specific age group represents that this age group has the narrowest gap with Vicuna in values. An increasing monoticity indicates a closer alignment towards younger groups, vice versa.
Figure 14: Alignment rank of LLMs over different age groups in Argentina and Brazil. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
Figure 15: Alignment rank of LLMs over different age groups in Ethiopia and Nigeria. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
Figure 16: Alignment rank of LLMs over different age groups in Gemany and Great Britain. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
Figure 17: Alignment rank of LLMs over different age groups in Indonesia and Malaysia. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.